# Closed-form approximations of the two-sample Pearson Bayes factor

### v3

@@ Abstract
Word count: 3,415
A common scenario in applied statistical inference involves comparing the means of two independent samples (e.g., a treatment group and a control group). This can be done using hypothesis testing (or more broadly, model comparison), where two hypotheses about potential differences between the two underlying population means $\mu_{1},\mu_{2}$ are compared after observing data. Classically, hypothesis tests work by first assuming a null hypothesis $\mathcal{H}_{0}$, and then calculating a test statistic that can be used to index the probability of obtaining some sample of observed data under the null hypothesis. If this probability is small, the common logic is to reject the null hypothesis $\mathcal{H}_{0}$ in favor of some alternative hypothesis $\mathcal{H}_{1}$. The problem tackled in the present paper is twofold. First, we consider an alternative method for this type of inference based on Bayes factors Kass and Raftery, (1995); Faulkenberry, (2025) and provide a novel, but simple method for their computation. Second, we compare this method with a classic approximation based on the Bayesian information criterion (BIC) and show that our new method outperforms the BIC method by a significant margin.
@@
Despite the popularity of this approach to testing for differences in sample means, there have been many recent criticisms against their use, and more generally, against null hypothesis significance testing Wagenmakers, (2007). Additionally, the American Statistical Association has recently recommended against the use of $p$-values and significance testing for scientific inference Wasserstein and Lazar, (2016). One alternative that has been recommended is a Bayesian approach, which works by considering the relative predictive adequacy of $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ against some observed data $\bm{y}$. This relative predictive adequacy is indexed by computing the Bayes factor Kass and Raftery, (1995), which indexes the extent to which the observed data $\bm{y}$ is more likely under one hypothesis – say $\mathcal{H}_{1}$ – compared to the other hypothesis – say $\mathcal{H}_{0}$. That is,
@@
In general, computing Bayes factors can be quite difficult, particularly because computing the marginal likelihoods involves nontrivial integration. One popular approach adopted in recent years is the BIC approximation Kass and Raftery, (1995); Wagenmakers, (2007); Masson, (2011), which works by constructing a second-order Taylor approximation of the log marginal likelihood about the posterior mode for $\theta_{i}$. Kass and Raftery Kass and Raftery, (1995) show that this results in the approximation
@@
This approximation uses the Schwarz Bayesian Information Criterion (BIC) Schwarz, (1978), which is computed as $\operatorname{BIC}(\mathcal{H}_{i})=-2\ln L_{i}+k_{i}\ln n$, where $k_{i}$ is the number of parameters in $\mathcal{H}_{i}$, $n$ is the number of data observations, and $L_{i}$ is the value of the likelihood function for model $\mathcal{H}_{i}$ with the maximum likelihood estimate taken as its argument. As typically used (e.g., our two-sample hypothesis testing problem), $n$ is taken as the number of distinct observations across all experimental groups (i.e., $n=2N$) and $k_{1}=k_{0}+1$. This last equality represents the fact that $\mathcal{H}_{1}$ has an additional parameter indexing the experimental group that $\mathcal{H}_{0}$ does not have. From here, it is easy to derive an approximation for the Bayes factor:
@@
Compared to computing marginal likelihoods by integrating the prior-weighted likelihoods, the BIC approximation requires only that we know the BIC values for models $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$. Generally, this requires that we have “raw” data available. Recent work by Faulkenberry Faulkenberry, (2018) improved the BIC method by making the Bayes factor computation accessible with only summary statistics:
@@
Despite the simplicity of computation offered by the BIC method, the fact remains that it is based on a large sample approximation, leaving it with limited utility in a small sample context. It is also limited for use in a teaching context, where sample sizes for hand-worked examples tend to be small. Recent work by Wang and Sun Wang and Sun, (2014) has provided another approach for calculating Bayes factors that can prove to be fruitful. In their work, Wang and Sun place a random effects linear model on the data
@@
Wang and Sun applied a method of Garcia-Donato and Sun García-Donato and Sun, (2007), who considered a proper prior on the ratio of variance components $\tau=\sigma^{2}_{a}/\sigma^{2}$ under $\mathcal{H}_{1}$. Under this prior specification, Garcia-Donato and Sun showed that the Bayes factor could be computed as
@@
where $\mathcal{B}(x,y)=\int_{0}^{1}t^{x-1}(1-t)^{y-1}dt$ is the standard Beta function. Wang and Liu Wang and Liu, (2016) further restricted $\pi^{PT}$ to one parameter $\alpha\in[-1,-\frac{1}{2}]$ by taking $\kappa=r$ and $\beta=\frac{n-p}{2}-\alpha-2$. A plot of this prior can be seen in Figure 1; in this figure, we assume $p=2$ groups and $r=20$ replicates per condition, giving a total of $n=40$ observations, thus setting $\kappa=20$ and $\beta=\frac{40-2}{2}-\alpha-2$, where $\alpha$ takes specific values between -1 and -1/2. Further, Figure 1 shows the effect of varying $\alpha$ on the prior distribution for $\tau$. As we can see, as $\alpha$ decreases from -1/2 to -1, $\tau$ becomes more dispersed and less peaked around the mode. This places more prior mass on larger treatment effects than we would see for values of $\alpha$ closer to $\frac{1}{2}$.
@@
With the above prior specification, Wang and colleagues Wang and Sun, (2014); Wang and Liu, (2016) proved that the Bayes factor of Garcia-Donato and Sun simplifies to an analytic expression without integral representation. The resulting Pearson Bayes factor Faulkenberry, 2020b allows one to calculate the Bayes factor for $\mathcal{H}_{1}$ over $\mathcal{H}_{0}$ without the need for integration:
@@
The Pearson Bayes factor is computable using only the summary statistics from the $t$-test (i.e., the $t$-score and the degrees of freedom $\nu$). It also includes a parameter $\alpha$ which allows the analyst to tune the scale of the prior distribution on effect sizes. Wang and Liu Wang and Liu, (2016) recommend a default setting of $\alpha=-1/2$, as the asymptotic tail behavior of the prior becomes a Cauchy distribution as $\tau$ increases, mirroring the prior structure recommended by Jeffreys Jeffreys, (1961). Following this recommendation, the Pearson Bayes factor simplifies to:
@@ 1 Approximating quotients of Gamma functions
Direct computation of these Gamma functions requires calculus (or more practically, numerical routines in a scientific programming language). Thus, the goal in this paper is to find closed-form approximations of this quotient that can be carried out using only basic algebraic operations. To this end, I have developed three such approximations – one that follows directly from a classical asymptotic formula of Wendel Wendel, (1948), one that derives directly from the classical Stirling formula Jameson, (2015), and finally, one that follows from an “improved” approximation of Frame Frame, (1949).
@@ 1.1 Wendel’s asymptotic formula
In his brief paper, Wendel Wendel, (1948) showed that for all real numbers $a$ and $x$,
@@ 1.2 Stirling’s formula
Another approach to approximating $C_{\nu}$ comes from applying Stirling’s formula Jameson, (2015). Historically, Stirling’s formula arose as a way to approximate the factorial function for the positive integers; i.e.,
@@ 1.3 Frame’s quotient formula
The final approach I will explore in this paper is derived from a method of Frame Frame, (1949), who proposed the following approximation to the quotient of two nearby values of the Gamma function:
@@ 2 Example computations
For illustration, let us now apply these three approximation methods to a concrete example. Consider the following summary data from Borota et al. Borota et al., (2014), who observed that with a sample of $n=73$ participants, those who received 200 mg of caffeine performed significantly better on a test of object memory compared to a control group of participants who received a placebo, $t(71)=2.0$, $p=0.049$. Borota and colleagues claimed this result as evidence that caffeine enhances memory consolidation.
@@
The approximation error we incur by using the Wendel asymptotic formula for approximating $C_{\nu}$ is small, resulting in an underestimate of 0.64460.6469=0.00230.6446-0.6469=-0.0023, a relative error magnitude of 0.36%. For comparison, consider the error that results from using the BIC method Kass and Raftery, (1995); Wagenmakers, (2007); Masson, (2011), a popular method for approximating Bayes factors direclty from summary statistics. Faulkenberry Faulkenberry, (2018) showed that the BIC Bayes factor can be computed directly as follows:
@@
Next, we will now apply the approximation based on Stirling’s formula to the Borota et al. Borota et al., (2014) summary statistics. This yields
@@
Finally, we will apply the approximation based on Frame’s formula to the Borota et al. Borota et al., (2014) summary statistics. This yields
@@ 3 Comparing the three approximations
    Using the stored values of $t$ and $\nu$, compute the BIC Bayes factor using the method of Faulkenberry Faulkenberry, (2018) and compute the Pearson Bayes factor using Equation 2, where $C_{\nu}$ is calculated two different ways:
@@ 4 Conclusion
In this paper, I have presented three new closed-form approximations of the two-sample Pearson Bayes factor. These techniques allow the user to compute reasonably accurate approximations for Bayes factors in two-sample designs without the need for computing the Gamma function. As such, these computations may be performed using nothing more than a simple scientific calculator, making them a very attractive option for users who wish to compute Bayes factors directly from summary statistics in two-sample designs. Though the formulas vary in complexity, even the simplest formula based on Wendel’s (1948) asymptotic formula produces Bayes factor approximations with average percent error dropping below 1% for reasonably small sample sizes. As all three are asymptotic methods, their relative error will decrease with increasing sample sizes. This is a much better approach to approximating Bayes factors compared to the often-used BIC approximation Kass and Raftery, (1995); Wagenmakers, (2007); Masson, (2011); Faulkenberry, (2018); Faulkenberry, 2020a ; Faulkenberry, (2019). The approximations presented here retain the spirit of the BIC Bayes factor (e.g., ease of use and ability to compute using only summary statistics), but as demonstrated, they provide a much better level of accuracy. One potential criticism of this new approach is that the presented approximations depend on a particular choice of prior (the Pearson Type VI prior). However, the same is true for the BIC approximation, which also assumes an underlying prior distribution (the unit information prior) Kass and Raftery, (1995). Given that the Bayes factors based on this Pearson Type VI prior show good performance against other well-known Bayes factor techniques Faulkenberry, 2020a , the approximations presented in this paper are the ideal tool for easily computing evidential value in two-sample designs.

### v4

@@ Abstract
Word count: 3,126
A common scenario in applied statistical inference involves comparing the means of two independent samples (e.g., a treatment group and a control group). This can be done using hypothesis testing (or more broadly, model comparison), where two hypotheses about potential differences between the two underlying population means $\mu_{1},\mu_{2}$ are compared after observing data. Classically, hypothesis tests work by first assuming a null hypothesis $\mathcal{H}_{0}$, and then calculating a test statistic that can be used to index the probability of obtaining some sample of observed data under the null hypothesis. If this probability is small, the common logic is to reject the null hypothesis $\mathcal{H}_{0}$ in favor of some alternative hypothesis $\mathcal{H}_{1}$. The problem tackled in the present paper is twofold. First, we consider an alternative method for this type of inference based on Bayes factors (Kass and Raftery,, 1995; Faulkenberry,, 2025) and provide a novel, but simple method for their computation. Second, we compare this method with a classic approximation based on the Bayesian information criterion (BIC) and show that our new method outperforms the BIC method by a significant margin.
@@
Despite the popularity of this approach to testing for differences in sample means, there have been many recent criticisms against their use, and more generally, against null hypothesis significance testing (Wagenmakers,, 2007). Additionally, the American Statistical Association has recently recommended against the use of $p$-values and significance testing for scientific inference (Wasserstein and Lazar,, 2016). One alternative that has been recommended is a Bayesian approach, which works by considering the relative predictive adequacy of $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ against some observed data $\bm{y}$. This relative predictive adequacy is indexed by computing the Bayes factor (Kass and Raftery,, 1995), which indexes the extent to which the observed data $\bm{y}$ is more likely under one hypothesis – say $\mathcal{H}_{1}$ – compared to the other hypothesis – say $\mathcal{H}_{0}$. That is,
@@
In general, computing Bayes factors can be quite difficult, particularly because computing the marginal likelihoods involves nontrivial integration. One popular approach adopted in recent years is the BIC approximation (Kass and Raftery,, 1995; Wagenmakers,, 2007; Masson,, 2011), which works by constructing a second-order Taylor approximation of the log marginal likelihood about the posterior mode for $\theta_{i}$. Kass and Raftery, (1995) show that this results in the approximation
@@
This approximation uses the Schwarz Bayesian Information Criterion (BIC) (Schwarz,, 1978), which is computed as $\operatorname{BIC}(\mathcal{H}_{i})=-2\ln L_{i}+k_{i}\ln n$, where $k_{i}$ is the number of parameters in $\mathcal{H}_{i}$, $n$ is the number of data observations, and $L_{i}$ is the value of the likelihood function for model $\mathcal{H}_{i}$ with the maximum likelihood estimate taken as its argument. As typically used (e.g., our two-sample hypothesis testing problem), $n$ is taken as the number of distinct observations across all experimental groups (i.e., $n=2N$) and $k_{1}=k_{0}+1$. This last equality represents the fact that $\mathcal{H}_{1}$ has an additional parameter indexing the experimental group that $\mathcal{H}_{0}$ does not have. From here, it is easy to derive an approximation for the Bayes factor:
@@
Compared to computing marginal likelihoods by integrating the prior-weighted likelihoods, the BIC approximation requires only that we know the BIC values for models $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$. Generally, this requires that we have “raw” data available. Recent work by Faulkenberry, (2018) improved the BIC method by making the Bayes factor computation accessible with only summary statistics:
@@
Despite the simplicity of computation offered by the BIC method, the fact remains that it is based on a large sample approximation, leaving it with limited utility in a small sample context. It is also limited for use in a teaching context, where sample sizes for hand-worked examples tend to be small. Recent work by Wang and Sun, (2014) has provided another approach for calculating Bayes factors that can prove to be fruitful. In their work, Wang and Sun place a random effects linear model on the data
@@
Wang and Sun applied a method of García-Donato and Sun, (2007), who considered a proper prior on the ratio of variance components $\tau=\sigma^{2}_{a}/\sigma^{2}$ under $\mathcal{H}_{1}$. Under this prior specification, Garcia-Donato and Sun showed that the Bayes factor could be computed as
@@
where $\mathcal{B}(x,y)=\int_{0}^{1}t^{x-1}(1-t)^{y-1}dt$ is the standard Beta function. Wang and Liu, (2016) further restricted $\pi^{PT}$ to one parameter $\alpha\in[-1,-\frac{1}{2}]$ by taking $\kappa=r$ and $\beta=\frac{n-p}{2}-\alpha-2$. A plot of this prior can be seen in Figure 1; in this figure, we assume $p=2$ groups and $r=20$ replicates per condition, giving a total of $n=40$ observations, thus setting $\kappa=20$ and $\beta=\frac{40-2}{2}-\alpha-2$, where $\alpha$ takes specific values between -1 and -1/2. Further, Figure 1 shows the effect of varying $\alpha$ on the prior distribution for $\tau$. As we can see, as $\alpha$ decreases from -1/2 to -1, $\tau$ becomes more dispersed and less peaked around the mode. This places more prior mass on larger treatment effects than we would see for values of $\alpha$ closer to $\frac{1}{2}$.
@@
With the above prior specification, Wang and colleagues (Wang and Sun,, 2014; Wang and Liu,, 2016) proved that the Bayes factor of Garcia-Donato and Sun simplifies to an analytic expression without integral representation. The resulting Pearson Bayes factor (Faulkenberry, 2020b, ) allows one to calculate the Bayes factor for $\mathcal{H}_{1}$ over $\mathcal{H}_{0}$ without the need for integration:
@@
The Pearson Bayes factor is computable using only the summary statistics from the $t$-test (i.e., the $t$-score and the degrees of freedom $\nu$). It also includes a parameter $\alpha$ which allows the analyst to tune the scale of the prior distribution on effect sizes. Wang and Liu, (2016) recommend a default setting of $\alpha=-1/2$, as the asymptotic tail behavior of the prior becomes a Cauchy distribution as $\tau$ increases, mirroring the prior structure recommended by Jeffreys, (1961). Following this recommendation, the Pearson Bayes factor simplifies to:
@@ 1 Approximating quotients of Gamma functions
Direct computation of these Gamma functions requires calculus (or more practically, numerical routines in a scientific programming language). Thus, the goal in this paper is to find closed-form approximations of this quotient that can be carried out using only basic algebraic operations. To this end, I have developed three such approximations – one that follows directly from a classical asymptotic formula of Wendel, (1948), one that derives directly from the classical Stirling formula (Jameson,, 2015), and finally, one that follows from an “improved” approximation of Frame, (1949).
@@ 1.1 Wendel’s asymptotic formula
In his brief paper, Wendel, (1948) showed that for all real numbers $a$ and $x$,
@@ 1.2 Stirling’s formula
Another approach to approximating $C_{\nu}$ comes from applying Stirling’s formula (Jameson,, 2015). Historically, Stirling’s formula arose as a way to approximate the factorial function for the positive integers; i.e.,
@@ 1.3 Frame’s quotient formula
The final approach I will explore in this paper is derived from a method of Frame, (1949), who proposed the following approximation to the quotient of two nearby values of the Gamma function:
@@ 2 Example computations
For illustration, let us now apply these three approximation methods to a concrete example. Consider the following summary data from Borota et al., (2014), who observed that with a sample of $n=73$ participants, those who received 200 mg of caffeine performed significantly better on a test of object memory compared to a control group of participants who received a placebo, $t(71)=2.0$, $p=0.049$. Borota and colleagues claimed this result as evidence that caffeine enhances memory consolidation.
@@
The approximation error we incur by using the Wendel asymptotic formula for approximating $C_{\nu}$ is small, resulting in an underestimate of 0.64460.6469=0.00230.6446-0.6469=-0.0023, a relative error magnitude of 0.36%. For comparison, consider the error that results from using the BIC method (Kass and Raftery,, 1995; Wagenmakers,, 2007; Masson,, 2011), a popular method for approximating Bayes factors direclty from summary statistics. Faulkenberry, (2018) showed that the BIC Bayes factor can be computed directly as follows:
@@
Next, we will now apply the approximation based on Stirling’s formula to the Borota et al., (2014) summary statistics. This yields
@@
Finally, we will apply the approximation based on Frame’s formula to the Borota et al., (2014) summary statistics. This yields
@@ 3 Comparing the three approximations
    Using the stored values of $t$ and $\nu$, compute the BIC Bayes factor using the method of Faulkenberry, (2018) and compute the Pearson Bayes factor using Equation 2, where $C_{\nu}$ is calculated two different ways:
@@ 4 Conclusion
In this paper, I have presented three new closed-form approximations of the two-sample Pearson Bayes factor. These techniques allow the user to compute reasonably accurate approximations for Bayes factors in two-sample designs without the need for computing the Gamma function. As such, these computations may be performed using nothing more than a simple scientific calculator, making them a very attractive option for users who wish to compute Bayes factors directly from summary statistics in two-sample designs. Though the formulas vary in complexity, even the simplest formula based on Wendel’s (1948) asymptotic formula produces Bayes factor approximations with average percent error dropping below 1% for reasonably small sample sizes. As all three are asymptotic methods, their relative error will decrease with increasing sample sizes. This is a much better approach to approximating Bayes factors compared to the often-used BIC approximation (Kass and Raftery,, 1995; Wagenmakers,, 2007; Masson,, 2011; Faulkenberry,, 2018; Faulkenberry, 2020a, ; Faulkenberry,, 2019). The approximations presented here retain the spirit of the BIC Bayes factor (e.g., ease of use and ability to compute using only summary statistics), but as demonstrated, they provide a much better level of accuracy. One potential criticism of this new approach is that the presented approximations depend on a particular choice of prior (the Pearson Type VI prior). However, the same is true for the BIC approximation, which also assumes an underlying prior distribution (the unit information prior) (Kass and Raftery,, 1995). Given that the Bayes factors based on this Pearson Type VI prior show good performance against other well-known Bayes factor techniques (Faulkenberry, 2020a, ), the approximations presented in this paper are the ideal tool for easily computing evidential value in two-sample designs.